LASER SPECTROSCOPIC STUDY OF THE SiAr VAN DER WAALS COMPLEX

CHONG TAO, \underline{ALEXEY TESLJA}, and PAUL J. DAGDIGIAN, Department of Chemistry, The Johns Hopkins University, Baltimore, Md 21218-2685; SULE ATAHAN and MILLARD H. ALEXANDER, Department of Chemistry and Biochemistry, The University of Maryland, College Park, MD 20742-2021.

Laser fluorescence excitation spectra of the SiAr van der Waals complex, in the vicinity of the Si ^{3}D^{o} ← ^{3}P atomic resonance transition near 220.7 nm are reported. At low resolution, a single excited state (v′,0) progression of bands terminating in a dissociation continuum is observed. Several weaker bands associated with many of these strong bands are found in scans at higher resolution. A transition to an excited ^{3}Σ^{−} state which correlates with the excited Si(^{3}D^{o})+Ar asymptote was assigned, and a rotational and vibrational analysis of the observed bands was carried out. The dissociation energies of the Ω = 0^{+} components of the ground X^{3}Σ^{−} and excited ^{3}Σ^{−} states were determined [D′′_{0} = 178.8 ± 0.4 and D′_{0} = 122.5 ± 0.4 cm^{−1}]. Ab initio calculations of the SiAr X^{3}Σ^{−} and A^{3}Π electronic states correlating with the groundstate Si(3s^{2}3p^{2} ^{3}P) + Ar asymptote were also carried out. The potential energy curves of the definite-Ω states were computed and used to estimate the dissociation energy, rotational constant, and phenomenological spin-spin interaction in the X^{3}Σ^{−} state. These parameters were found to be in resonable agreement with the experimental determinations.